OpenStudy (anonymous):

solve for variable -x + -58 = -20 I got 78 please correct me.

OpenStudy (anonymous):

the correct answer is x=-38

hero (hero):

@mathdummy4 , you should provide steps, not answers. If you want to prove that what you say is true, provide the steps. Providing an answer only doesn't prove anything and it is against CoC.

OpenStudy (anonymous):

i'm sorry and also i did my math wrong mode....follow hero's advice its correct...again my appologies

hero (hero):

-x - 58 = -20 -(x + 58) = -20 x + 58 = -20/-1 x + 58 = 20 x = 20 - 58 x = ? Yes, I made a mistake, but @mathdummy4 , it would have been better for you to just post the correct solution than to just post the answer. Either way OS doesn't like that.

OpenStudy (anonymous):

well like i've already said...i apologize for not following OS CoC...im still fairly new to the sight and have not thoroughly read over the rules...i didn't mean to call anyone out...